# Relative_Constructibility_via_Restricted_ASMs.bib

### v1

@@ 2.7 Transfinite Runs
The symbols in $\sigma(\mathfrak{M})$ have default interpretations following $\mathfrak{M}$. For all $\ulcorner X\urcorner\in\sigma(\mathfrak{M})$, the default interpretation of $\ulcorner R^{\ulcorner X\urcorner}\urcorner$ is the empty set. The two extra constant symbols $\ulcorner c_{0}\urcorner$ and $\ulcorner c_{1}\urcorner$ default to 00 and $1$ respectively.
At every step of a computation of $\mathfrak{M}^{*}$, it checks if “$c_{0}=c_{1}$”. If the answer is no, it copies the previous $\mathfrak{M}$ portion data to the $R^{\ulcorner X\urcorner}$s within the fibers denoted by $c_{0}$ (which is supposed to count the steps in every short terminating run), simulates the next step of its $\mathfrak{M}$ portion, and increments both $c_{0}$ and $c_{1}$. Otherwise, the moral implication is that a limit stage has been reached, whence it first verifies if $Y$ is functional on the current $R^{\ulcorner X\urcorner}$s (note that the $R^{\ulcorner X\urcorner}$s might contain gibberish in states not reachable from any initial state). If so, it computes point-wise limits of the $R^{\ulcorner X\urcorner}$s according to $Y$, copies the resulting limit state information to its $\mathfrak{M}$ portion, and increments just $c_{1}$. Otherwise, it terminates. We can argue by induction that $\mathfrak{M}^{*}$ always correctly computes the states of a short terminating run of $\mathfrak{M}$, primarily because the preservation of ordinal limits by the simple limit formulas in $Z$ allows both
@@
In conclusion, the $\mathfrak{M}$ portion of $\mathfrak{M}^{*}$ simulates (with some latency) the entirety of each short terminating run of $\mathfrak{M}$. As said portion contains both the input and output tapes of $\mathfrak{M}^{*}$, (3) holds. ∎
@@
Basically, we are mandating the taking of limit inferiors point-wise wherever possible to be the standard method by which we compute limit states of transfinite runs. This is not a real restriction, for we have established from Propositions 2.37 and 2.41 that our new standard is among the most powerful ways to take limits of an RASM’s states, while still maintaining absoluteness and the spirit of local definability.
Until now, we have only considered “short” transfinite runs, that is, runs that take less steps than the length of the input/output tape. What about longer runs? After all, the ability for Turing machines to run forever is the reason they can “output” non-computable computably enumerable sets. Allowing a Turing machine to run transfinitely (e.g. in [ittm]) allows it to explicitly decide these sets. Through such means, oracle computation becomes more straightforward: the oracle can be incorporated into the input, and the result of the machine being fed that oracle can be directly written on the output tape. We want to have an analogous concept of “long” runs on RASMs. It turns out that removing the length constraint on terminating runs works just fine;
@@
Let $\mathfrak{M}$ be an RASM. A terminating run of $\mathfrak{M}$ with input $A$ and output $B$ is a function $Sq$ from some successor ordinal $\delta$ into $S(\mathfrak{M})$ such that
@@
$\text{terminating run of }\mathfrak{M}\text{ with input }A\text{ and output }B\}$
@@
$\text{terminating run of }\mathfrak{M}\text{ with input }A\text{ and output }B\}$
@@
For a limit ordinal $\kappa$ and two sets of ordinals $A$ and $B$, that $\mathrm{Runs}_{\kappa}(A,B)$ is a set and not a proper class stems from the fact that by our definition of terminating runs, $Sq$ must be an injection into $S(\mathfrak{M})$ whenever $(\mathfrak{M},Sq)\in\mathrm{Runs}_{\kappa}(A,B)$. The aforementioned fact also means that all terminating runs of an RASM $\mathfrak{M}$ must be of length $<(2^{\kappa(\mathfrak{M})})^{+}$.
@@ 3 Computability and Comparisons
In this section, we derive from our restricted abstract state machines, various notions of generalised (relative) computability. We then compare these notions with existing ones.
@@ 3.1 Relative Computability
  • for some RASM $\mathfrak{M}$ with $\kappa(\mathfrak{M})=\kappa$, there is a terminating run of $\mathfrak{M}$ with input $A$ and output $B$.
@@
    $A$ is $\kappa$-computable iff $A\leq^{C}_{\kappa}\emptyset$.
@@
    $A$ is computable iff $A$ is $\kappa$-computable for some limit ordinal $\kappa$.
@@
Let $\mathfrak{M}_{1}$ witness $B\leq^{C}_{\kappa}C$ and $\mathfrak{M}_{2}$ witness $A\leq^{C}_{\kappa}B$. We shall construct an RASM $\mathfrak{M}^{*}$ with $\kappa(\mathfrak{M}^{*})=\kappa$ that essentially simulates $\mathfrak{M}_{1}$ followed by $\mathfrak{M}_{2}$, so that $\mathfrak{M}^{*}$ witnesses $A\leq^{C}_{\kappa}C$. Set $\sigma(\mathfrak{M}^{*})$ to be a disjoint union of $\sigma(\mathfrak{M}_{1})$ and $\sigma(\mathfrak{M}_{2})$, and
@@
It makes sense that a machine is more powerful when endowed with more space. This might not quite be the case for RASMs, but we can get close to this sort of monotonicity with a mild assumption.
@@
Let $\mathfrak{M}$ be an RASM witnessing $A\leq^{C}_{\kappa}B$. Define an RASM $\mathfrak{M}^{\prime}$ with $\kappa(\mathfrak{M}^{\prime})=\kappa^{\prime}$ and $\sigma(\mathfrak{M}^{\prime})=\sigma(\mathfrak{M})$. Since $\kappa$ is definable over $(\kappa^{\prime};\in)$, we can relativise both $\phi^{d}_{\mathfrak{M}}$ and $\phi^{\tau}_{\mathfrak{M}}$ to $\kappa^{\prime}$ so that if $\ulcorner X\urcorner\in\sigma(\mathfrak{M})$ is an $n$-ary function symbol and $\alpha_{1},\dots,\alpha_{n}\in\kappa^{\prime}$, then
@@ 3.2 Further Machine Modifications
for all limit ordinals $\kappa$ and $B\subset\kappa$. An ideal relative computability relation uniformly definable in $\kappa$ should have its predecessor sets grow along with $\kappa$. Towards such a relation, we shall introduce an augmentation to RASMs.
@@
An RASM with parameters (henceforth RASMP)
@@
is defined like an RASM, except constants in $\sigma(\mathfrak{M})$ are allowed to evaluate to ($(\{\alpha\},1)$ under $\vartheta_{\mathfrak{M}}$ for any $\alpha\in\kappa(\mathfrak{M})$. In this case, we call $\alpha$ a parameter of $\mathfrak{M}$.
@@
Fix any limit ordinal $\kappa$. There are $|\kappa|$ many RASMPs $\mathfrak{M}$ satisfying $\kappa(\mathfrak{M})=\kappa$. Also, this set of RASMP specifications is absolute for transitive models of $\mathsf{ZFC}$ containing $\kappa$ (cf. Remark 2.26).
The definitions and properties so far associated with RASMs adapt perfectly to RASMPs. We will use the adapted definitions in the context of RASMPs without further clarifications.
@@
  • for some RASMP $\mathfrak{M}$ with $\kappa(\mathfrak{M})=\kappa$, there is a terminating run of $\mathfrak{M}$ with input $A$ and output $B$.
@@
    $A$ is $\kappa$-computable with parameters iff $A\leq^{P}_{\kappa}\emptyset$.
@@
    $A$ is computable with parameters iff $A$ is $\kappa$-computable with parameters for some limit ordinal $\kappa$.
@@
$\text{terminating run of }\mathfrak{M}\text{ with input }A\text{ and output }B\}$
@@
$\text{terminating run of }\mathfrak{M}\text{ with input }A\text{ and output }B\}$
@@
Adapting Proposition 2.28 to RASMPs, in conjunction with Remark 3.7, the absoluteness of $\kappa^{<\omega}$, and the absoluteness of the definition of the $\mathrm{RunsP}_{\kappa}(A,B)$s, gives us that $\leq^{P}_{\kappa}$ is absolute for transitive models of $\mathsf{ZFC}$ containing $\kappa$. As a result, like $\leq^{C}$, $\leq^{P}$ is absolute for transitive models of $\mathsf{ZFC}$ with the same ordinals.
@@
As there is no way to formulate $\phi^{d}_{\mathfrak{M}^{\prime}}$ such that $\ulcorner c\urcorner$ occurs in it, we choose $\phi^{d}_{\mathfrak{M}^{\prime}}$ to be any universally valid sentence over $\sigma(\mathfrak{M^{\prime}})$. For example, setting all return values to 00 works. Define $\tau_{\mathfrak{M}^{\prime}}$ as follows. If “$d=0$” in the current state, make use of $c$ to set up the $\sigma(\mathfrak{M})$ portion of the next state according to the relativised version of $\phi^{d}_{\mathfrak{M}}$, then increment $d$. Otherwise, do nothing to $d$ and modify the $\sigma(\mathfrak{M})$ portion of the next state according to $\phi^{\tau}_{\mathfrak{M}}$ relativised to $c$. The $\mathfrak{M}^{\prime}$ defined thus witnesses $A\leq^{P}_{\kappa^{\prime}}B$. ∎
@@
Proposition 3.12, the transitivity of $\leq^{P}$, and the nice relationship between limit ordinals $\kappa$ and the sizes of $\leq^{P}_{\kappa}$ predecessor sets, all point to $\leq^{P}$, $\leq^{P}_{\kappa}$ being the superior and more valid notion of relative computability. In fact, they also point to an RASMP being a better model of generalised algorithm than an RASM. For this reason, the rest of this subsection will focus on $\leq^{P}$, $\leq^{P}_{\kappa}$ and RASMPs.
We are now able to solidify our case for using point-wise limit inferiors, whenever they exist, to compute limit states of terminating runs which are not short. Let us first generalise the concepts of ordinal limits preservation and simple limit states (Definitions 2.35 and 2.36) to RASMPs and sequences of arbitrary lengths.
@@
Let $\mathfrak{M}$ be an RASMP. Then $\mathfrak{M}$ is strongly closed under a set $Y$ iff for each $n$-ary function symbol $\ulcorner X\urcorner\in\sigma(\mathfrak{M})$ there is a simple limit formula $\phi_{lim}^{\ulcorner X\urcorner}$ for $(\ulcorner X\urcorner,\mathfrak{M})$, such that
@@
Let $\mathfrak{M}$ be an RASMP. Then $\mathfrak{M}$ is practically closed under a set $Y$ iff for each $n$-ary function symbol $\ulcorner X\urcorner\in\sigma(\mathfrak{M})$ there is a simple limit formula $\phi_{lim}^{\ulcorner X\urcorner}$ for $(\ulcorner X\urcorner,\mathfrak{M})$, such that
@@
  • substituting “RASMP” for “RASM” and
  • removing the requirement for $\delta$ to be $<\kappa(\mathfrak{M})$.
@@
Proposition 2.37 holds with “RASMP” substituted for “RASM” and the requirement for $\delta$ to be $<\kappa(\mathfrak{M})$ removed.
@@
Any formula that works for all RASMs $\mathfrak{M}$ with $\ulcorner X\urcorner\in\sigma(\mathfrak{M})$ in the context of proving Proposition 2.37 (such a formula exists by the last sentence of the proof of Proposition 2.37) also works here. ∎
@@
Suppose $\mathfrak{M}$ is an RASMP practically closed under $Y$. Then for some signature $\sigma^{\prime}$, as long as $\mathfrak{M}^{\prime}$ is an RASMP satisfying
@@
and $Z$ is a set of simple limit formulas preserving ordinal limits with $\mathfrak{M}^{\prime}$ strongly closed under $Z$, there is an RASM $\mathfrak{M}^{*}$ such that
  1. (1)
@@
Let us try to follow the proof of Proposition 2.41 with $\mathfrak{M}^{\dagger}:=\mathcal{G}(\mathfrak{M},(2^{\kappa(\mathfrak{M})})^{+})$ specified in place of the arbitrary “$\mathfrak{M}$” given by said proposition. Note that
  • Remark 2.44 remains true with “$\mathrm{RunsP}_{\kappa}(A,B)$” substituted for “$\mathrm{Runs}_{\kappa}(A,B)$” and “RASMP” substituted for “RASM”, and
  • (3) holds with $\mathfrak{M}^{\dagger}$ in place of $\mathfrak{M}^{*}$.
@@
If classical recursion or computability theory is about the study of subsets of $\omega$ through Turing reducibility and degrees, then one can envision a natural extension of this study aiming at analogous notions of reducibility and degrees for subsets of ordinals larger than $\omega$. Indeed, there have been a number of successful extensions, lumped under the subject of higher recursion theory. The theories of alpha recursion (or $\alpha$-recursion) and $E$-recursion are the most developed among them, but both recursion theories in their typical presentations are built upon schemata akin to Gödel’s model of real computation (see [takeuti] and [normann] respectively).
Emerging later into the scene is the theory of ordinal computability, started by Koepke in [koepke1]. A comparison with $\alpha$-recursion theory is given in [koepke2]. We will drawn on parts of this comparative study to give an equivalent definition of $\leq^{P}$.
Before that, we need a brief introduction to select set-theoretic concepts, for set theory is unavoidable in the study of sets beyond $\omega$.

### v2

@@ 2.7 Transfinite Runs
The symbols in $\sigma(\mathfrak{M})$ have default interpretations following $\mathfrak{M}$. For all $\ulcorner X\urcorner\in\sigma(\mathfrak{M})$, the default interpretation of $\ulcorner R^{\ulcorner X\urcorner}\urcorner$ is the empty set. The two extra constant symbols $\ulcorner c_{0}\urcorner$ and $\ulcorner c_{1}\urcorner$ default to 0 and $1$ respectively.
At every step of a computation of $\mathfrak{M}^{*}$, it checks if $c_{0}=c_{1}$”. If the answer is no, it copies the previous $\mathfrak{M}$ portion data to the $R^{\ulcorner X\urcorner}$s within the fibers denoted by $c_{0}$ (which is supposed to count the steps in every short terminating run), simulates the next step of its $\mathfrak{M}$ portion, and increments both $c_{0}$ and $c_{1}$. Otherwise, the moral implication is that a limit stage has been reached, whence it first verifies if $Y$ is functional on the current $R^{\ulcorner X\urcorner}$s (note that the $R^{\ulcorner X\urcorner}$s might contain gibberish in states not reachable from any initial state). If so, it computes point-wise limits of the $R^{\ulcorner X\urcorner}$s according to $Y$, copies the resulting limit state information to its $\mathfrak{M}$ portion, and increments just $c_{1}$. Otherwise, it terminates. We can argue by induction that $\mathfrak{M}^{*}$ always correctly computes the states of a short terminating run of $\mathfrak{M}$, primarily because the preservation of ordinal limits by the simple limit formulas in $Z$ allows both
@@
In conclusion, the $\mathfrak{M}$ portion of $\mathfrak{M}^{*}$ simulates (with some latency) the entirety of each short terminating run of $\mathfrak{M}$. As said portion contains both the input and output tapes of $\mathfrak{M}^{*}$, (c) holds. ∎
@@
Basically, we are mandating the taking of limit inferiors point-wise wherever possible to be the standard method by which we compute limit states of transfinite runs. This is not a real restriction, for we have established from Propositions 2.37 and 2.41 that our new standard is among the most powerful ways to take limits of a GSeqA’s states, while still maintaining absoluteness and the spirit of local definability.
Until now, we have only considered “short” transfinite runs, that is, runs that take less steps than the length of the input/output tape. What about longer runs? After all, the ability for Turing machines to run forever is the reason they can “output” non-computable computably enumerable sets. Allowing a Turing machine to run transfinitely (e.g. in [8]) allows it to explicitly decide these sets. Through such means, oracle computation becomes more straightforward: the oracle can be incorporated into the input, and the result of the machine being fed that oracle can be directly written on the output tape. We want to have an analogous concept of “long” runs on GSeqAs. It turns out that removing the length constraint on terminating runs works just fine;
@@
Let $\mathfrak{M}$ be a GSeqA. A terminating run of $\mathfrak{M}$ with input $A$ and output $B$ is a function $Sq$ from some successor ordinal $\delta$ into $S(\mathfrak{M})$ such that
  1. (a)
@@
[…]
@@
[…]
@@
For a limit ordinal $\kappa$ and two sets of ordinals $A$ and $B$, that $\mathrm{Runs}_{\kappa}(A,B)$ is a set and not a proper class stems from the fact that by our definition of terminating runs, $Sq$ must be an injection into $S(\mathfrak{M})$ whenever $(\mathfrak{M},Sq)\in\mathrm{Runs}_{\kappa}(A,B)$. The aforementioned fact also means that all terminating runs of a GSeqA $\mathfrak{M}$ must be of length $(2^{\kappa(\mathfrak{M})})^{+}$.
@@ 3 Computability and Comparisons
In this section, we derive from our generalised sequential algorithms, various notions of generalised (relative) computability. We then compare these notions with existing ones.
@@ 3.1 Relative Computability
  2. (b)
    for some GSeqA $\mathfrak{M}$ with $\kappa(\mathfrak{M})=\kappa$, there is a terminating run of $\mathfrak{M}$ with input $A$ and output $B$.
@@
    AA is $\kappa$-computable iff $A\leq^{C}_{\kappa}\emptyset$.
@@
    AA is computable iff $A$ is $\kappa$-computable for some limit ordinal $\kappa$.
@@
Let $\mathfrak{M}_{1}$ witness $B\leq^{C}_{\kappa}C$ and $\mathfrak{M}_{2}$ witness $A\leq^{C}_{\kappa}B$. We shall construct a GSeqA $\mathfrak{M}^{*}$ with $\kappa(\mathfrak{M}^{*})=\kappa$ that essentially simulates $\mathfrak{M}_{1}$ followed by $\mathfrak{M}_{2}$, so that $\mathfrak{M}^{*}$ witnesses $A\leq^{C}_{\kappa}C$. Set $\sigma(\mathfrak{M}^{*})$ to be a disjoint union of $\sigma(\mathfrak{M}_{1})$ and $\sigma(\mathfrak{M}_{2})$, and
@@
It makes sense that a machine is more powerful when endowed with more space. This might not quite be the case for GSeqAs, but we can get close to this sort of monotonicity with a mild assumption.
@@
Let $\mathfrak{M}$ be a GSeqA witnessing $A\leq^{C}_{\kappa}B$. Define a GSeqA $\mathfrak{M}^{\prime}$ with $\kappa(\mathfrak{M}^{\prime})=\kappa^{\prime}$ and $\sigma(\mathfrak{M}^{\prime})=\sigma(\mathfrak{M})$. Since $\kappa$ is definable over $(\kappa^{\prime};\in)$, we can relativise both $\phi^{d}_{\mathfrak{M}}$ and $\phi^{\tau}_{\mathfrak{M}}$ to $\kappa^{\prime}$ so that if $\ulcorner X\urcorner\in\sigma(\mathfrak{M})$ is an $n$-ary function symbol and $\alpha_{1},\dots,\alpha_{n}\in\kappa^{\prime}$, then
  1. (1)
@@ 3.2 Further Modifications
for all limit ordinals $\kappa$ and $B\subset\kappa$ (which follows from Remark 2.26). An ideal relative computability relation uniformly definable in $\kappa$ should have its predecessor sets grow along with $\kappa$. Towards such a relation, we shall introduce an augmentation to GSeqAs.
@@
An GSeqA with parameters (henceforth GSeqAP)
@@
is defined like a GSeqA, except constants in $\sigma(\mathfrak{M})$ are allowed to evaluate to $(\{\alpha\},1)$ under $\vartheta_{\mathfrak{M}}$ for any $\alpha\in\kappa(\mathfrak{M})$. In this case, we call $\alpha$ a parameter of $\mathfrak{M}$.
@@
Fix any limit ordinal $\kappa$. There are $|\kappa|$ many GSeqAPs $\mathfrak{M}$ satisfying $\kappa(\mathfrak{M})=\kappa$. Also, this set of GSeqAP specifications is absolute for transitive models of $\mathsf{ZFC}$ containing $\kappa$ (cf. Remark 2.26).
The definitions and properties so far associated with GSeqAs adapt perfectly to GSeqAPs. We will use the adapted definitions in the context of GSeqAPs without further clarifications.
@@
  2. (b)
    for some GSeqAP $\mathfrak{M}$ with $\kappa(\mathfrak{M})=\kappa$, there is a terminating run of $\mathfrak{M}$ with input $A$ and output $B$.
@@
    AA is $\kappa$-computable with parameters iff $A\leq^{P}_{\kappa}\emptyset$.
@@
    AA is computable with parameters iff $A$ is $\kappa$-computable with parameters for some limit ordinal $\kappa$.
@@
[…]
@@
[…]
@@
Adapting Proposition 2.28 to GSeqAPs, in conjunction with Remark 3.7, the absoluteness of $\kappa^{<\omega}$, and the absoluteness of the definition of the $\mathrm{RunsP}_{\kappa}(A,B)$s, gives us that $\leq^{P}_{\kappa}$ is absolute for transitive models of $\mathsf{ZFC}$ containing $\kappa$. As a result, like $\leq^{C}$, $\leq^{P}$ is absolute for transitive models of $\mathsf{ZFC}$ with the same ordinals.
@@
As there is no way to formulate $\phi^{d}_{\mathfrak{M}^{\prime}}$ such that $\ulcorner c\urcorner$ occurs in it, we choose $\phi^{d}_{\mathfrak{M}^{\prime}}$ to be any universally valid sentence over $\sigma(\mathfrak{M^{\prime}})$. For example, setting all return values to 0 works. Define $\tau_{\mathfrak{M}^{\prime}}$ as follows. If “d=0d=0” in the current state, make use of $c$ to set up the $\sigma(\mathfrak{M})$ portion of the next state according to the relativised version of $\phi^{d}_{\mathfrak{M}}$, then increment $d$. Otherwise, do nothing to $d$ and modify the $\sigma(\mathfrak{M})$ portion of the next state according to $\phi^{\tau}_{\mathfrak{M}}$ relativised to $c$. The $\mathfrak{M}^{\prime}$ defined thus witnesses $A\leq^{P}_{\kappa^{\prime}}B$. ∎
@@
Proposition 3.12, the transitivity of $\leq^{P}$, and the nice relationship between limit ordinals $\kappa$ and the sizes of $\leq^{P}_{\kappa}$ predecessor sets, all point to $\leq^{P}$, $\leq^{P}_{\kappa}$ being the superior and more valid notion of relative computability. In fact, they also point to a GSeqAP being a better model of generalised algorithm than a GSeqA. For this reason, the rest of this subsection will focus on $\leq^{P}$, $\leq^{P}_{\kappa}$ and GSeqAPs.
We are now able to solidify our case for using point-wise limit inferiors, whenever they exist, to compute limit states of terminating runs which are not short. Let us first generalise the concepts of ordinal limits preservation and simple limit states (Definitions 2.35 and 2.36) to GSeqAPs and sequences of arbitrary lengths.
@@
Let $\mathfrak{M}$ be a GSeqAP. Then $\mathfrak{M}$ is strongly closed under a set $Y$ iff for each $n$-ary function symbol $\ulcorner X\urcorner\in\sigma(\mathfrak{M})$ there is a simple limit formula $\phi_{lim}^{\ulcorner X\urcorner}$ for $(\ulcorner X\urcorner,\mathfrak{M})$, such that
@@
Let $\mathfrak{M}$ be a GSeqAP. Then $\mathfrak{M}$ is practically closed under a set $Y$ iff for each $n$-ary function symbol $\ulcorner X\urcorner\in\sigma(\mathfrak{M})$ there is a simple limit formula $\phi_{lim}^{\ulcorner X\urcorner}$ for $(\ulcorner X\urcorner,\mathfrak{M})$, such that
@@
  • substituting “GSeqAP” for “GSeqA” and
  • removing the requirement for $\delta$ to be $\kappa(\mathfrak{M})$.
@@
Proposition 2.37 holds with “GSeqAP” substituted for “GSeqA” and the requirement for $\delta$ to be $\kappa(\mathfrak{M})$ removed.
@@
Any formula that works for all GSeqAs $\mathfrak{M}$ with $\ulcorner X\urcorner\in\sigma(\mathfrak{M})$ in the context of proving Proposition 2.37 (such a formula exists by the last sentence of the proof of Proposition 2.37) also works here. ∎
@@
Suppose $\mathfrak{M}$ is a GSeqAP practically closed under $Y$. Then for some signature $\sigma^{\prime}$, as long as $\mathfrak{M}^{\prime}$ is a GSeqAP satisfying
@@
and $Z$ is a set of simple limit formulas preserving ordinal limits with $\mathfrak{M}^{\prime}$ strongly closed under $Z$, there is a GSeqAP $\mathfrak{M}^{*}$ such that
  1. (a)
@@
Let us try to follow the proof of Proposition 2.41 with $\mathfrak{M}^{\dagger}:=\mathcal{G}(\mathfrak{M},(2^{\kappa(\mathfrak{M})})^{+})$ specified in place of the arbitrary $\mathfrak{M}$” given by said proposition. Note that
  • Remark 2.44 remains true with $\mathrm{RunsP}_{\kappa}(A,B)$” substituted for $\mathrm{Runs}_{\kappa}(A,B)$” and “GSeqAP” substituted for “GSeqA”, and
  • (c) holds with $\mathfrak{M}^{\dagger}$ in place of $\mathfrak{M}^{*}$.
@@
If classical recursion or computability theory is about the study of subsets of $\omega$ through Turing reducibility and degrees, then one can envision a natural extension of this study aiming at analogous notions of reducibility and degrees for subsets of ordinals larger than $\omega$. Indeed, there have been a number of successful extensions, lumped under the subject of higher recursion theory. The theories of alpha recursion (or $\alpha$-recursion) and $E$-recursion are the most developed among them, but both recursion theories in their typical presentations are built upon schemata akin to Gödel’s model of real computation (see [1] and [3] respectively).
Emerging later into the scene is the theory of ordinal computability, started by Koepke in [13]. A comparison with $\alpha$-recursion theory is given in [14]. We will drawn on parts of this comparative study to give an equivalent definition of $\leq^{P}$.